OpenStudy (anonymous):

simplify the imaginary number √-12

OpenStudy (michele_laino):

hint: we can write this: \[\Large \sqrt { - 12} = \sqrt { - 1 \times {2^2} \times 3} = ...?\]

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

can you just give the answer this one time so it will refresh my memory

OpenStudy (michele_laino):

hint: \[\Large \sqrt { - 12} = \sqrt { - 1 \times {2^2} \times 3} = \sqrt { - 1} \times 2 \times \sqrt 3 = ...?\]

OpenStudy (michele_laino):

please keep in mind that: \[i = \sqrt { - 1} \]

OpenStudy (anonymous):

im sorry i dont know what the answer is going to look like

OpenStudy (anonymous):

thx anyway

OpenStudy (michele_laino):

ok! Here is your answer: \[\Large \sqrt { - 12} = \sqrt { - 1 \times {2^2} \times 3} = \sqrt { - 1} \times 2 \times \sqrt 3 = 2i\sqrt 3 \]